OpenStudy (anonymous):

*********medals and fans***********Courtney sold 50 cantaloupes at the farmers' market and had 27 left. Which equation could be used to find x, the number of cantaloupes Courtney had originally?

OpenStudy (anonymous):

x - 27 = 50 x - 50 = -27 x + 50 = 27 x + 27 = -50

OpenStudy (anonymous):

x - 27 = 50

OpenStudy (anonymous):

x = no. of cantaloupes x - 27 = 50